A shop has 5 different printers but there is space for only 3 printers on the display shelf. How many arrangements are possible i. Suppose you have 4 different flags. How many different signals could you make using (i) 2 flags (ii) 2 or 3 flags ii. iii. How many arrangements of the letters of the word B E G I N are there if (i) 3 letters are used (ii) all of the letters are used A quiz team of four is chosen from a group of 15 students. In how many ways could the team be chosen? iv. If there are eight girls and seven boys in a class, in how many ways could a group be chosen so that there are two boys and two girls in the group? v. A school committee consists of six girls and four boys. A social sub-committee consisting of four students is to be formed. In how many ways could the group be chosen if there are to be more girls than boys in the group?


1
Expert's answer
2021-05-31T16:26:06-0400

1.

"N=C^3_5\cdot3!=\frac{5!}{3!2!}\cdot3!=60"


2.

i)

"N=2C^2_4=\frac{4!}{2!2!}\cdot2=12"


ii)

"N=2C^2_4+C^3_4\cdot3!=12+\frac{4!}{3!}\cdot3!=36"


3.

i)

"N=C^3_5\cdot3!=\frac{5!}{3!2!}\cdot3!=60"


ii)

"N=5!=120"


4.

"N=C^4_{15}=\frac{15!}{11!4!}=1365"


5.

"N=C^2_7\cdot C^2_8=\frac{7!}{5!2!}\cdot\frac{8!}{6!2!}=21\cdot28=588"


6.

Total number of ways=number of groups with 3 girls + number of groups with 4 girls.

number of groups with 3 girls:

"N_1=C^3_6\cdot C^1_4=\frac{6!}{3!3!}\cdot\frac{4!}{3!}=20\cdot4=80"

number of groups with 4 girls:

"N_2=C^4_6=\frac{6!}{4!2!}=15"

Total number of ways:

"N=N_1+N_2=80+15=95"